# Efficiently Serving LLM Reasoning Programs with Certaindex

### v1

@@ 3 Certaindex Based Resource Allocation
Large language models (LLMs) have an inherent ability to "know when they know", meaning they can self-assess their confidence in their answers. This ability, discovered in prior work [21], allows LLMs to indicate their certainty while generating tokens. In this section, we develop certaindex, a measure of this confidence, as a proxy for reasoning progress: high certainty suggests the LLM is nearing a final answer or directly indicates a lower absolute compute needed to reach a final answer, whether correct or not. Our goal is to measure and track certaindex during inference, enabling dynamic resource allocation, prioritization of complex queries, scaling back simpler queries, and terminating unpromising queries.
@@ 3.1 Measuring Certaindex
Certaindex in typical reasoning algorithms. Reasoning algorithms like SC, MCTS, and Rebase expand multiple reasoning paths to derive aggregated final answers (§2.1). To quantify LLM’s certainty among these paths, we employ semantic entropy [22], which is derived from empirical entroy. Given a question $P$, $n$ reasoning paths are generated and clustered into $m$ groups based on their answers: $C_{1},C_{2},...,C_{m}$, where $|C_{i}|$ denotes the number of paths in answer group $C_{i}$. The empirical entropy is calculated as: $\mathcal{H}=-\sum_{i=1}^{m}\frac{|C_{i}|}{n}\log\frac{|C_{i}|}{n}$.
@@ 3.2 Effectiveness of Certaindex
Pareto-frontier allocation. Ideally, a more refined allocation, depicted by the fitted green curve, dynamically determines the maximum inference compute allocation for each certaindex value. In addition to the thresholding-based allocation, which only impacts queries with certaindex above $t$, this allocation caps resources for all queries at any certaindex while preserving accuracy. We discuss the comparison of these two allocation strategies for specific dataset/model/reasoning algorithm combinations in §6.4.
3.3 Comparing Certaindex with Other Signals
@@
Using the (SC, GSM8K, Llama3.1-8B-instruct) setup, we evaluated certaindex alongside these metrics. Four metrics were tested: certaindex’s entropy measure $\mathcal{H}$, mean output length [13], mean normalized log probability [31], and their linear combinations. As shown in Fig. 5, $\mathcal{H}$ achieved the strongest correlation with ground truth compute requirements (Pearson Correlation of 0.68, Kendall’s Tau of 0.61), outperforming other metrics and matching complex combinations. These results confirm that certaindex is a simple yet effective proxy for estimating inference computational demands, offering robust performance across tasks and models. Our end-to-end token-to-accuracy evaluations in §6.5 compare different signals for resource allocation, confirming certaindex’s superior performance.
@@
4 Characterizing the Relationship Between Certaindex and Detection Steps
Fig. 4 demonstrates the relationship between certaindex values and program steps across different detection points in a single run, using SC on the GSM8K dataset with Llama3.1 8B Instruct model. The figure consists of six subplots, each representing a different detection step ranging from 5 to 30, indicated by "Detect @knob" values.
@@ 4 Characterizing the Relationship Between Certaindex and Detection Steps
5 Dynasor: System Design
We present Dynasor, an end-to-end LLM serving system that optimizes reasoning programs using certaindex. Dynasor provides a simple interface supporting diverse reasoning algorithms and integrates resource allocation and request scheduling for efficient LLM serving.
Fig. 6 illustrates the architecture of Dynasor, which consists of three main components: (a) a Reasoning Program Abstraction (Program) offering a unified interface for reasoning algorithms, (b) an Application Runtime, which dynamically allocates resources based on certaindex, and (c) a System Runtime, managing request-level scheduling on the backend.
@@ 5 Dynasor: System Design
5.1 Reasoning Program Abstraction
@@
5.2 Application Runtime
5.2.1 Certaindex-Based Intra-Program Scheduler
@@
Resource allocation policy. A resource allocation policy determines compute allocation (e.g., branches in SC and iterations in MCTS) for each program based on its certaindex. §3.2 discussed two certaindex-based allocation policies: simple thresholding and pareto-frontier, both implemented in the system. Additional alternatives are analyzed in §6.4. Developers can easily configure the intra-program scheduler to use other certaindex-based policies.
5.2.2 Profiler-Guided Policy Calibration
We explain how to determine (or calibrate) a resource allocation policy, focusing on a certaindex threshold policy as our example. Determining an effective resource allocation policy is critical for reasoning program performance (§6.4). Program submitters often lack insight into runtime patterns, and program characteristics may shift due to algorithmic changes or data distribution shifts.
@@ 5.2.2 Profiler-Guided Policy Calibration
5.3 System Runtime
5.3.1 Program-Aware Inter-Program Scheduler
@@
5.3.2 Other System Runtime Components
@@ 5.4 System Implementation
Using Dynasor, we adapt various reasoning algorithms (§2.1) to the Program interface, including their custom certaindex implementations. Our system comprises approximately 3,000 lines of Python code, covering the Program interface, application runtime, and system runtime.
@@
6 Evaluation
Figure 8: Token-to-accuracy metric on batch processing workloads. Mean performance and std (error bars) of 10 runs are reported.
In this section, we evaluate Dynasor across four reasoning algorithms (SC [48] , Rebase [51], MCTS [16, 12], and ICoT) on diverse datasets [17, 7, 34, 19] in both offline and online settings. Our experiments demonstrate that Dynasor consistently outperforms state-of-the-art LLM serving systems. Specifically, for offline workloads (§6.2), Dynasor reduces token usage by up to $52.2\%$ while maintaining accuracy requirements. In online serving (§6.3), Dynasor handles up to $\times3$ higher request rates and meets 4.7×4.7$\times$4.7 × stricter SLOs, meeting latency constraints for over $90\%$ of requests (P90) compared to leading serving engines like SGLang [57] and ParrotServe [29]. Additionally, we perform ablation studies to analyze the contributions of different system components.
6.1 Experimental Setup
General Settings. Tables 2 and 3 summarize our offline and online evaluation workloads, respectively. Table 1 lists the certaindex hyperparameters for each workload, including thresholds (tuned on test sets) and detection steps (detect @knob). In our main experiments, we use a threshold-based policy (§3.2 and §5.2.1): we scale knob and detect certaindex at the detection step (detect @knob), terminating the program if its certaindex values meet all threshold conditions. All experiments run on a GPU cluster equipped with A100 (80GB) GPUs. GPU usage varies by method: Rebase and MCTS use two GPUs to separately serve their LLM and reward models, while SC requires only one GPU for request processing. For final answer selection, SC uses simple majority voting, Rebase applies weighted majority voting, and MCTS selects the best path using the reward model via best-of-n.

### v2

@@ Appendix B Certaindex Based Resource Allocation
Large language models (LLMs) have an inherent ability to “know when they know”, meaning they can self-assess their confidence in their answers. This ability, discovered in prior work such as  [22], allows LLMs to indicate their certainty while generating tokens. In this section, We introduce certaindex, a measure of this confidence, as a proxy for reasoning progress: high certainty suggests the LLM is nearing a final answer or directly indicates a lower absolute compute needed to reach a final answer, whether correct or not. Our goal is to measure and track certaindex during inference, enabling dynamic resource allocation, prioritization of complex queries, scaling back simpler queries, and terminating unpromising queries. Some of the contents will overlap with § 3.1, as we want to provide a more complete narrative of the certaindex definition.
@@ B.1 Measuring Certaindex
Certaindex in typical reasoning algorithms. Reasoning algorithms like SC, MCTS, and Rebase expand multiple reasoning paths to derive aggregated final answers (Appendix A). To quantify LLM’s certainty among these paths, we employ semantic entropy [23], which is derived from empirical entroy. Given a question $P$, $n$ reasoning paths are generated and clustered into $m$ groups based on their answers: $C_{1},C_{2},...,C_{m}$, where $|C_{i}|$ denotes the number of paths in answer group $C_{i}$. The empirical entropy is calculated as: $\mathcal{H}=-\sum_{i=1}^{m}\frac{|C_{i}|}{n}\log\frac{|C_{i}|}{n}$.
@@ B.2 Effectiveness of Certaindex
Pareto-frontier allocation. While thresholding offers a coarse-grained control, a more nuanced approach is a Pareto-frontier allocation. Instead of a binary stop/go decision based on a single early checkpoint, this policy aims to assign a dynamically tailored computational budget to each query based on its evolving certaindex throughout the reasoning process. The underlying principle is to leverage the observed correlation to inform a more continuous resource allocation. Rather than simply “fitting a curve”, we establish an empirically-derived relationship that maps different certaindex values to an estimated optimal remaining token budget for that level of certainty. This mapping (conceptually represented by the green curve in Figure 12) reflects a trade-off: for higher certaindex values, the optimal additional budget is small, while for lower certaindex values, more computation might be warranted, up to a certain point. This allows for a dynamic allocation scheme where the maximum additional tokens for a query are continuously adjusted or capped based on its current certaindex. Queries exhibiting high certaindex (signaling proximity to a stable solution) are allocated fewer future tokens, while those still showing lower certaindex (indicating ongoing exploration or difficulty) might be allowed more, guided by this learned relationship. This strategy seeks to optimize resource use more globally across all queries, distributing the computational budget more effectively, as further elaborated in our “smart scheduler” design (§G.4).
B.3 Comparing Certaindex with Other Signals
@@
Using the (SC, GSM8K, Llama3.1-8B-instruct) setup, we evaluated certaindex alongside these metrics. Four metrics were tested: certaindex’s entropy measure $\mathcal{H}$, mean output length [41], mean normalized log probability [25], and their linear combinations. As shown in Fig. 13, $\mathcal{H}$ achieved the strongest correlation with ground truth compute requirements (Pearson Correlation of 0.68, Kendall’s Tau of 0.61), outperforming other metrics and matching complex combinations. These results confirm that certaindex is a simple yet effective proxy for estimating inference computational demands, offering robust performance across tasks and models. Our end-to-end token-to-accuracy evaluations in Appendix G compare different signals for resource allocation, confirming certaindex’s superior performance.
@@
Appendix C Characterizing the Relationship Between Certaindex and Detection Steps
Fig. 14 demonstrates the relationship between certaindex values and program steps across different detection points in a single run, using SC on the GSM8K dataset with Llama3.1 8B Instruct model. The figure consists of six subplots, each representing a different detection step ranging from 5 to 30, indicated by “Detect @knob” values.
@@ Appendix C Characterizing the Relationship Between Certaindex and Detection Steps
Appendix D Dynasor: System Design
Figure 15 illustrates Dynasor’s three-component architecture: (a) a Reasoning Program Abstraction (Program) offering a standardized interface for diverse reasoning algorithms, (b) an Application Runtime that dynamically allocates computational resources based on CertaIndex measurements, and (c) a System Runtime managing request-level scheduling on the backend infrastructure.
This architecture enables Dynasor to adapt computation dynamically to each query’s difficulty level, allocating resources proportionally to reasoning complexity rather than using static, one-size-fits-all token budgets.
@@ Appendix D Dynasor: System Design
D.1 Reasoning Program Abstraction
@@
D.2 Application Runtime
D.2.1 Certaindex-Based Intra-Program Scheduler
@@
Resource allocation policy. A resource allocation policy determines compute allocation (e.g., branches in SC and iterations in MCTS) for each program based on its certaindex. §B.2 discussed two certaindex-based allocation policies: simple thresholding and pareto-frontier, both implemented in the system. Additional alternatives are analyzed in §G.4. Developers can easily configure the intra-program scheduler to use other certaindex-based policies.
D.2.2 Profiler-Guided Policy Calibration
We explain how to determine (or calibrate) a resource allocation policy, focusing on a certaindex threshold policy as our example. Determining an effective resource allocation policy is critical for reasoning program performance (§G.4). Program submitters often lack insight into runtime patterns, and program characteristics may shift due to algorithmic changes or data distribution shifts.
@@ D.2.2 Profiler-Guided Policy Calibration
D.3 System Runtime
D.3.1 Program-Aware Inter-Program Scheduler
@@
D.3.2 Other System Runtime Components
@@ D.4 System Implementation
Using Dynasor, we adapt various reasoning algorithms (Appendix A) to the Program interface, including their custom certaindex implementations. Our system comprises approximately 1.5k lines of Python code, covering the Program interface, application runtime, and system runtime. Notably, the core system runtime only comprises around $\sim 500$ lines of code change, and the changes are modular and non-invasive, making it a very clean implementation into the core serving system logic.
@@
Appendix E Proof of Stopping Criteria for CoT and Probing
Given the input prompt $x$, we can stop the reasoning chain $Y_{1},\dots,Y_{t},Y_{t+1}$ at time $t$ if the following holds for any $T>t$.
$P_{t}(Y_{t+1}|x,Y_{1},\dots,Y_{t})\stackrel{{\scriptstyle\mathrm{t}}}{{=}}P_{t+1}(Y_{t+2}|x,Y_{1},\dots,Y_{t+1})\stackrel{{\scriptstyle\mathrm{t+1}}}{{=}}\dots\stackrel{{\scriptstyle\mathrm{T}}}{{=}}P_{T}(Y_{T+1}|x,Y_{1},\dots,Y_{T})=P_{*}(Y|x).$ (2)
The above definition means that adding new steps in the reasoning chain does not change the distribution of the generated answer anymore. As a result, the reasoning chain reaches the stationary distribution, $P_{*}(Y|x)$, which is the desired output given the prompt $x$ in the first place.
We assume that the language model will eventually converge in the chain of thought, if the answer distribution remains the same for long enough.
Formally it is expressed as follows.
Assumption 1.
If equality $P_{t}(Y_{t+1}|x,Y_{1},\dots,Y_{t})=P_{t+i}(Y_{t+i+1}|x,Y_{1},\dots,Y_{t+i})$ holds for any $1\leq i\leq k^{*}$,
then $P_{t}(Y_{t+1}|x,Y_{1},\dots,Y_{t})=P_{T}(Y_{T+1}|x,Y_{1},\dots,Y_{T})=P_{*}$ holds for any $T>t$. Further assume that $k^{*}=\mathcal{O}(M)$, where $M$ is the number of distinct output groups.
Based on the above assumption, we propose the test criterion in Definition 2.
Definition 1.
Consider $n$ samples, each corresponds to $P_{t},\dots,P_{t+n}$. Denote the mixture distribution of samples $i+1,\dots,i+k\in[1,\dots,n]$ to be $\bar{P}_{i}^{i+k}=\frac{1}{k}\sum_{j=1}^{k}P_{i+j}$.
Let the smallest $t$ where Eq. (2) holds be $t^{*}$. Then for any $i,j\geq t^{*}$, $\bar{P}_{i}^{i+k}=\bar{P}_{j}^{j+k}=P_{*}$. On the other hand, if $i<t^{*}$, then there exists $j>i$, so that $\bar{P}_{i}^{i+k}\neq\bar{P}_{j}^{j+k}$. This intuition is formally expressed in Lemma 2 below. We therefore propose to test the difference among $\bar{P}_{i}^{i+k}$, $\bar{P}_{i+1}^{i+k+1},\dots,\bar{P}_{i+k}^{i+2k}$.
Lemma 2.
If $\mathrm{TV}(\bar{P}_{i}^{i+k},\bar{P}_{i+j}^{i+j+k})=0$, for any $1\leq j\leq k$, and if $\mathrm{TV}(\bar{P}_{i}^{i+k-1},\bar{P}_{i+j}^{i+j+k-1})=0$, for any $11\leq j\leq k-$ 1, then Eq. (2) holds for any $t\geq i$, and $T<2k-1$.
Proof.
We first have from the definition of the TV distance that
$\left\|\bar{P}_{i}^{i+k}-\bar{P}_{i+j}^{i+j+k}\right\|_{1}=0.$
For a general $1\leq j\leq k$, we have from the definition of $\bar{P}$ that
$\frac{1}{k}\left\|\sum_{l=1}^{j}\left(P_{i+l}-P_{i+l+k}\right)\right\|_{1}=\left\|\bar{P}_{i}^{i+k}-\bar{P}_{i+j}^{i+j+k}\right\|_{1}=0.$
Therefore, $\sum_{l=1}^{j}P_{i+l}=\sum_{l=1}^{j}P_{i+l+k}$.
Take $j=1$, then
$P_{i+1}=P_{i+1+k}$. Plugging into the case where $j=2$, then we have $P_{i+2}=P_{i+2+k}$. Expanding the recursion for all $1\leq j\leq k$, we obtain that
$P_{i+j}=P_{i+j+k},\quad\forall 1\leq j\leq k.$ (3)
Similarly from the assumption that $\mathrm{TV}\left(\bar{P}_{i}^{i+k-1},\bar{P}_{i+j}^{i+j+k-1}\right)=0$, for any $11\leq j\leq k-$ 1, we have that:
$0=\left\|\bar{P}_{i}^{i+k-1}-\bar{P}_{i+j}^{i+j+k-1}\right\|_{1}=\frac{1}{k-1}\left\|\sum_{l=1}^{j}\left(P_{i+l}-P_{i+l+k-1}\right)\right\|_{1},\quad\forall 1\leq j\leq k-1,$
and consequently that
$P_{i+j}=P_{i+j+k-1},\quad\forall 1\leq j\leq k-1.$ (4)
The equalities in Eqs. (3) and (4) form a graph over the vertices that are $P_{i+1},\dots,P_{i+2k-1}$, where the equalities are the edges. Note that all the nodes are edge connected by virtue of Eqs. (3) and (4). Therefore, we have that
$P_{j}=P_{l},\quad\forall i+1\leq j,l\leq 2k-1.$
We therefore define the approximate stopping criteria as follows.
Definition 2.
We define the $\epsilon$-accuracy of the stopping criteria:
$\mathrm{TV}(\bar{P}_{i}^{i+k},\bar{P}_{i+j}^{i+j+k})\leq\epsilon,\text{ for any }1\leq j\leq k;$
$\mathrm{TV}(\bar{P}_{i}^{i+k-1},\bar{P}_{i+j}^{i+j+k-1})\leq\epsilon,\text{ for any }1\leq j\leq k-1.$
We then ask: how many data samples $k$ do we need, to be able to test our stopping criteria up to $\epsilon$-accuracy? More concretely, we construct the estimator of $\bar{P}_{l}^{l+t}$, for $l=i+1,\dots,i+k$, and $t=k-1,k$ as follows. Assume we have $M$ disjoint groups of the outputs. Denote each group of answer as $C_{m}$, for $m=1,\dots,M$. Then we estimate the average probability mass function $\bar{P}_{l}^{l+t}$ by defining the Bernoulli random variable for each $C_{m}$: $Z^{C_{m}}_{\tau}=\mathbbm{1}\{Y_{\tau}\in C_{m}\}$, and $\hat{P}_{l}^{l+t}(S)=\frac{1}{t}\sum_{\tau=l+1}^{l+t}Z^{S}_{\tau}$.
The next lemma establishes that with $k=\widetilde{\Omega}\left(\frac{M+\log(1/\delta)}{\epsilon^{2}}\right)$, then $\mathrm{TV}(\bar{P}_{l}^{l+t},\hat{P}_{l}^{l+t})\leq\epsilon/3$, for all $l=i+1,\dots,i+k$, and $t=k-1,k$, with $1-\delta$ probability. Then we just have to test that
$\mathrm{TV}(\hat{P}_{i}^{i+k},\hat{P}_{i+j}^{i+j+k})\leq\epsilon/3,\text{ for any }1\leq j\leq k;$
$\mathrm{TV}(\hat{P}_{i}^{i+k-1},\hat{P}_{i+j}^{i+j+k-1})\leq\epsilon/3,\text{ for any }1\leq j\leq k-1.$
If the above inequalities stand with the established number of steps $k$, then by the triangle inequality of the total variation (TV) distance, we obtain the $\epsilon$-accuracy in the stopping criteria in Definition 2:
$\mathrm{TV}(\bar{P}_{i}^{i+k},\bar{P}_{i+j}^{i+j+k})\leq\mathrm{TV}(\bar{P}_{i}^{i+k},\hat{P}_{i}^{i+k})+\mathrm{TV}(\hat{P}_{i}^{i+k},\hat{P}_{i+j}^{i+j+k})+\mathrm{TV}(\hat{P}_{i+j}^{i+j+k},\bar{P}_{i+j}^{i+j+k})\leq\epsilon.$
Lemma 1.
If $k=\widetilde{\Omega}\left(\frac{M+\log(1/\delta)}{\epsilon^{2}}\right)$, then $\mathrm{TV}(\bar{P}_{l}^{l+t},\hat{P}_{l}^{l+t})\leq\epsilon/3$, for all $l=i+1,\dots,i+k$, and $t=k-1,k$, with $1-\delta$ probability.
Proof.
To prove convergence in the total variation (TV) distance, we first express it as follows. Let $S$ belong to the power set (denoted as $\exp^{C}$) of $\bigcup_{m=1}^{M}C_{m}$. Then
$\mathrm{TV}(\bar{P},\hat{P})=\sup_{S\in\exp^{C}}\left|\bar{P}(S)-\hat{P}(S)\right|.$
We then prove convergence for a fixed set $S$, and then apply union bound to $\exp^{C}$ to obtain the final result.
Denote a Bernoulli random variable $Z^{S}_{\tau}=\mathbbm{1}\{Y_{\tau}\in S\}$. Then $\hat{P}_{l}^{l+t}(S)=\frac{1}{t}\sum_{\tau=l+1}^{l+t}Z^{S}_{\tau}$. On the other hand, $\bar{P}_{l}^{l+t}(S)=\frac{1}{t}\sum_{\tau=l+1}^{l+t}P_{\tau}(Y_{\tau}\in S|x,Y_{1},\dots,Y_{\tau-1})$,
Note that $Y_{\tau}|x,Y_{1},\dots,Y_{\tau-1}\sim P_{\tau}(\cdot|x,Y_{1},\dots,Y_{\tau-1})$. Therefore,
$\mathbb{E}\left[Z^{S}_{\tau}|x,Y_{1},\dots,Y_{\tau-1}\right]=P_{\tau}(Y_{\tau}\in S|x,Y_{1},\dots,Y_{\tau-1}).$
This means that $Z^{S}_{\tau}-P_{\tau}(Y_{\tau}\in S|x,Y_{1},\dots,Y_{\tau-1})$ is a martingale difference sequence. We also know that $\left|Z^{S}_{\tau}-P_{\tau}(Y_{\tau}\in S|x,Y_{1},\dots,Y_{\tau-1})\right|\leq 1$. Therefore, by the Azuma-Hoeffding inequality, we have:
$\mathbb{P}\left(\left|\sum_{\tau=l+1}^{l+t}\left(Z^{S}_{\tau}-P_{\tau}(Y_{\tau}\in S|x,Y_{1},\dots,Y_{\tau-1})\right)\right|\geq\epsilon\right)\leq 2\exp\left(-\frac{\epsilon^{2}}{2t}\right),$
or
$\mathbb{P}\left(\left|\bar{P}_{l}^{l+t}(S)-\hat{P}_{l}^{l+t}(S)\right|\geq\epsilon\right)$ $=\mathbb{P}\left(\left|\frac{1}{t}\sum_{\tau=l+1}^{l+t}Z^{S}_{\tau}-\frac{1}{t}\sum_{\tau=l+1}^{l+t}P_{\tau}(Y_{\tau}\in S|x,Y_{1},\dots,Y_{\tau-1})\right|\geq\epsilon\right)$
$\leq 2\exp\left(-\frac{t\epsilon^{2}}{2}\right).$
Taking a union bound over $S\in\exp^{C}$, we have:
$\mathbb{P}\left(\mathrm{TV}(\bar{P}_{l}^{l+t},\hat{P}_{l}^{l+t})\geq\epsilon\right)$ $=\mathbb{P}\left(\sup_{S\in\exp^{C}}\left|\bar{P}_{l}^{l+t}(S)-\hat{P}_{l}^{l+t}(S)\right|\geq\epsilon\right)$
$\leq 2^{M}\mathbb{P}\left(\left|\bar{P}_{l}^{l+t}(S)-\hat{P}_{l}^{l+t}(S)\right|\geq\epsilon\right)\leq 2^{M+1}\exp\left(-\frac{t\epsilon^{2}}{2}\right).$
On top of that, taking a union bound over $l=i+1,\dots,i+k$, and $t=k-1,k$, we have that the probability of $\mathrm{TV}(\bar{P}_{l}^{l+t},\hat{P}_{l}^{l+t})\geq\epsilon$ for one of the combination of $l=i+1,\dots,i+k$, and $t=k-1,k$ is less than:
$\sum_{l=i+1}^{k}\sum_{t=k-1}^{k}\mathbb{P}\left(\mathrm{TV}(\bar{P}_{l}^{l+t},\hat{P}_{l}^{l+t})\geq\epsilon\right)\leq 2^{M+2}k\cdot\exp\left(-\frac{(k-1)\epsilon^{2}}{2}\right).$
In other words, whenever
$k=\widetilde{\Omega}\left(\frac{M+\log(1/\delta)}{\epsilon^{2}}\right),$
we have $\mathrm{TV}(\bar{P}_{l}^{l+t},\hat{P}_{l}^{l+t})\leq\epsilon/3$, for all $l=i+1,\dots,i+k$, and $t=k-1,k$, with $1-\delta$ probability. Here $\widetilde{\Omega}(\cdot)$ means that we omit logarithmic order terms in $M$ and $\epsilon$.
Appendix F Detailed Setups for Offline and Online Experiments
(Algo., Dataset, LLM) Reward Model # Samples Resource Cap
(SC, LCB, Llama3.1 8B) / 400 5,10,15,20,25,30
(SC, GSM8K, Llama3.1 8B) / 1000 5,10,15,20,25,30
(MCTS, ASDiv, Llama2 7B) Skywork 7B 300 3,7,10,15,20
(MCTS, GSM8K, Llama2 7B) Skywork 7B 300 3,7,10,15,20
(Rebase, MATH, Llemma 7B) Llemma 34b 500 16,32,64,128
(Rebase, GSM8K, Llemma 34B) Llemma 34b 500 16,32,64,128
Table 2: Online workload Configurations
Algorithm Dataset LLM Reward Model Base Deadline (s)
SC MATH Llama3.1 8B / 240
MCTS ASDiv Llama2 7B Skywork 7B 60
Rebase GSM8K Llemma 34B Llemma 34b 300